OpenStudy (mckenzieandjesus):

The waiter places a bowl of soup in front of Abbey. In a counterclockwise direction, she passes the soup to Kai who then passes it to Haifa. Which two rotations about the center of the table describe passing the soup?

OpenStudy (mckenzieandjesus):

first 80° and then 120° first 40° and then 60° first 60° and then 90° first 72° and then 108°

OpenStudy (mckenzieandjesus):

please help!

jimthompson5910 (jim_thompson5910):

How many people are at the table?

OpenStudy (mckenzieandjesus):

13

OpenStudy (mckenzieandjesus):

no 12

jimthompson5910 (jim_thompson5910):

close, 14 actually

jimthompson5910 (jim_thompson5910):

oops my bad

OpenStudy (mckenzieandjesus):

but u skipped numbers?

jimthompson5910 (jim_thompson5910):

I jumped from 10 to 12 for some strange reason

OpenStudy (mckenzieandjesus):

no 12

jimthompson5910 (jim_thompson5910):

sorry I see where I added the extra number

jimthompson5910 (jim_thompson5910):

ok yeah, 12

OpenStudy (mckenzieandjesus):

ok lol

jimthompson5910 (jim_thompson5910):

anyways, divide 360 degrees by 12 360/12 = 30

jimthompson5910 (jim_thompson5910):

going from person to person means that you go 30 degrees along the outer arc

jimthompson5910 (jim_thompson5910):

so going from Abbey to Kai means we go 2 spots (2 spots)*(30 degrees per spot) = 60 degree rotation

OpenStudy (mckenzieandjesus):

yeah but the next spot i over 180? if its 30 degree each spot

jimthompson5910 (jim_thompson5910):

going from Kai to Haifa means we move 3 spots (3 spots)*(30 degrees per spot) = 90 degree rotation

OpenStudy (mckenzieandjesus):

its 2 spots and the 5 spots?

OpenStudy (mckenzieandjesus):

ohh i see lol

jimthompson5910 (jim_thompson5910):

yeah it's starting from Kai on the second part (and not Abbey)

OpenStudy (mckenzieandjesus):

Okay do u know this one? Every _____ tessellates. pentagon octagon triangle decagon

jimthompson5910 (jim_thompson5910):

the answer is given at the top of this page http://gabrielattar.weebly.com/uploads/4/3/2/2/43226303/12.6_tessaletions_notes_.pdf where it says "Theorem 12-6"

OpenStudy (mckenzieandjesus):

yay i got it right then

jimthompson5910 (jim_thompson5910):

congrats on getting it right

OpenStudy (mckenzieandjesus):

A taxicab starts at (1, -2) on the grid and picks up a passenger. The taxicab travels a total of 16 blocks, and then drops the passenger off. Which composition of two translations could describe the taxicab’s final position? (x, y) > (x + 4, y - 3), and then (x, y) >(x - 9, y + 1). (x, y) > (x - 4, y + 3), and then (x, y) > (x + 6, y - 1). (x, y) > (x - 3, y + 4), and then (x, y) > (x + 8, y - 1). My answer (x, y) > (x - 3, y + 4), and then (x, y) > (x + 7, y - 1).

OpenStudy (mckenzieandjesus):

Thanks (:

jimthompson5910 (jim_thompson5910):

good, choice C is the only one where the total distance traveled is 16 blocks

OpenStudy (mckenzieandjesus):

So im right!?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (mckenzieandjesus):

Yes! (:

jimthompson5910 (jim_thompson5910):

Nice job

OpenStudy (mckenzieandjesus):

In the diagram, triangle ECD is a _____ of triangle ABC across line p.

OpenStudy (mckenzieandjesus):

translation glide reflection rotation reflection (my guess)

jimthompson5910 (jim_thompson5910):

let me think

OpenStudy (mckenzieandjesus):

ok

jimthompson5910 (jim_thompson5910):

this is what I get when I reflect ABC over line p

jimthompson5910 (jim_thompson5910):

the reflection does not line up with ECD so reflection doesn't work I couldn't get glide reflection to work either

jimthompson5910 (jim_thompson5910):

translation doesn't work either the only thing left is rotation. The only issue is that I'm not sure how they're rotating it. This problem is just odd

jimthompson5910 (jim_thompson5910):

if line p were along BC, then a glide reflection would work but that's not the case since line p is through EC

OpenStudy (mckenzieandjesus):

hmm

jimthompson5910 (jim_thompson5910):

I think the authors messed up

OpenStudy (mckenzieandjesus):

All i can guess is relfection

OpenStudy (mckenzieandjesus):

A blueprint for a fence has a scale of 1:40. The fence in the blueprint is 6 in. What is the length of the actual fence? 20 feet 240 feet (or this one) 20 inches (my answer) 72 feet

OpenStudy (mckenzieandjesus):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

"has a scale of 1:40" means that you multiply the smaller dimension (6 inches) by 40 to get the corresponding longer dimension then convert to feet

OpenStudy (mckenzieandjesus):

Yeah. 6*40=240 240/12 = 20 right?

jimthompson5910 (jim_thompson5910):

looks good

OpenStudy (mckenzieandjesus):

ok